OpenStudy (anonymous):

Which answer correct estimates the product of 3.09 × 304.87? A. about 9.0 B. about 9.9 C. about 90 D. about 900 Plz explane it to me

OpenStudy (gabbyalicorn):

To multiply 304.87 by 3.09, set them up as shown by arranging the two numbers in columns. We usually put the number with the smaller number of digits at the bottom. Firstly, work out 9 × 30487. As mentioned before, we can ignore the decimal point for now. The smaller figures in green are the tens figures carried over. Working from right to left: 9 × 7 = 63 Put the 3 in the appropriate column and carry the 6. 9 × 8 + 6(carried) = 78 Put the 8 in the appropriate column and carry the 7. 9 × 4 + 7(carried) = 43 Put the 3 in the appropriate column and carry the 4. 9 × 0 + 4(carried) = 4 9 × 3 = 27 Put the 7 in the appropriate column and carry the 2. We now work out 00 × 30487. This can be written as 0 × 30487 × 10. Because we are multiplying by 10, we place 1 zero to the right and then work out 0 × 30487. Since we are multiplying 30487 by 0, this line could be left out altogether. 0 × 7 = 0 0 × 8 = 0 0 × 4 = 0 0 × 0 = 0 0 × 3 = 0 Lastly, work out 300 × 30487. This can be written as 3 × 30487 × 100. Because we are multiplying by 100, we place 2 zeroes to the right and then work out 3 × 30487. 3 × 7 = 21 Put the 1 in the appropriate column and carry the 2. 3 × 8 + 2(carried) = 26 Put the 6 in the appropriate column and carry the 2. 3 × 4 + 2(carried) = 14 Put the 4 in the appropriate column and carry the 1. 3 × 0 + 1(carried) = 1 3 × 3 = 9 Finally the 3 rows are added together starting at the right. The smaller figures in green are the tens figures carried over. 0 + 0 + 3 = 3 0 + 0 + 8 = 8 1 + 0 + 3 = 4 6 + 0 + 4 = 10 Put the 0 in the appropriate column and carry the 1. 4 + 0 + 7 + 1(carried) = 12 Put the 2 in the appropriate column and carry the 1. 1 + 0 + 2 + 1(carried) = 4 9 = 9 How is the decimal point positioned? 304.87 has 2 digits after the decimal point and 3.09 also has 2 digits after the decimal point making a total of 4 digits after the decimal point. The final answer must also have 4 digits after the decimal point. So: 304.87 × 3.09 = 942.0483

OpenStudy (gabbyalicorn):

Forget that last screenshot

OpenStudy (gabbyalicorn):

Make sense?

OpenStudy (anonymous):

yea, so it would be D?

OpenStudy (gabbyalicorn):

Yes, because anything below 5 we round down.

OpenStudy (anonymous):

Thank you for you'r help :)

OpenStudy (gabbyalicorn):

You're absolutely welcome. :)